OpenStudy (anonymous):

What is lim (cos((pi/2)+h) - cos(pi/2))/h? h->0

ganeshie8 (ganeshie8):

Hint : Let \(f(x)=\cos(x)\) the given expression is equivalent to \(f'(\pi/2)\)

OpenStudy (anonymous):

\[\frac{ \sin(h)-0}{ h }\]

OpenStudy (anonymous):

\[\frac{ \sin(h) }{ h }\]

ganeshie8 (ganeshie8):

\[\large \frac{ \color{red}{-}\sin(h)-0}{ h }\] right ?

ganeshie8 (ganeshie8):

because \(\cos(\pi/2 + x) = - \sin(x)\)

OpenStudy (anonymous):

ohhh ok! Is that like a identity?

ganeshie8 (ganeshie8):

do you remember angle sum identity ? \(\cos(A+B)\)

OpenStudy (anonymous):

sort of, it sounds familiar

ganeshie8 (ganeshie8):

\(\cos(A+B) = \cos A\cos B - \sin A\sin B\)

ganeshie8 (ganeshie8):

use that to work out \(\cos(\pi/2+h)\)

OpenStudy (anonymous):

ohhhhhhh that makes so much sense now

OpenStudy (anonymous):

\[\frac{ -\sin(h) }{ h }\]

OpenStudy (anonymous):

So do I substitute 0 for h now?

ganeshie8 (ganeshie8):

substitute h=0 and see what you get

OpenStudy (anonymous):

Oh yea that doesn't work with 0 on the bottom oops!

ganeshie8 (ganeshie8):

right, try using the standard limits 

OpenStudy (anonymous):

So -sin(h)/h =-1?

ganeshie8 (ganeshie8):

Yes : \[\large \lim\limits_{h\to 0}~\dfrac{ -\sin(h) }{ h } =- \lim\limits_{h\to 0}~\dfrac{ \sin(h) }{ h } =-1 \]

OpenStudy (anonymous):

Ohh I understand it all now!!! Thank you so much for teaching me you're great! :)

ganeshie8 (ganeshie8):

np :)